OpenStudy (anonymous):

An airline had a policy of maintaining a certain ratio of the number of members of cabin crew to the number of passengers, on every flight. A certain flight had a six member cabin crew on board. The following equation represents this information, where x is the maximum number of people, including the crew, on that flight. x = 6 + 23/24 x How many passengers were on that flight? Answer 136 150 144 138

OpenStudy (anonymous):

Like any other equation, the first step is to get all the variable terms on one side and the constants on the other.

OpenStudy (anonymous):

so its.... 150?

OpenStudy (anonymous):

Subtract 23/24x from both sides. x = 6 + (23/24)x -(23/24)x - (23/24)x (1/24)x = 6 Multiply both sides by 24 x=6(24) = 144